OpenStudy (anonymous):

pls. help the volume of a rectangular parallelepiped is 162 cubic units. the three dimensions are in the ratio 1:2:3. find the total surface area

OpenStudy (anonymous):

V = (length) (width) (height) The dimensions are 1:2:3 so: V = 1 (length) * 2 (width) * 3 (height) Because the values don't matter and are all relative to one another we can assume they are all base x. V = 1 (x) * 2 (x) * 3 (x)

OpenStudy (anonymous):

Solve for x first, then find the surface area

OpenStudy (anonymous):

Twice the surface of the base down= 2(x (2x)) 4 x^2

OpenStudy (anonymous):

Lateral Area = perimeter of the base times 3x = ( x+ 2x + x + 2x)(3x)=18x^2 Total area = 4 x^2 + 18 x^2 = 22 x^2 Notice There is an equal sign missing between 2(x (2x)) and 4 x^2 It should have been 2(x (2x)) =4 x^2

OpenStudy (anonymous):

If you know x, you will be done. 162 = x (2x)(3x) =6 x^3 Solve it, you find x=3

OpenStudy (anonymous):

Total area 22 x^2 = 22(9) =198